OpenStudy (anonymous):

the area of a circle is 81pi? what is the diameter of the circle

terenzreignz (terenzreignz):

What's the formula for the area of a circle?

OpenStudy (anonymous):

Area = π × r2

terenzreignz (terenzreignz):

Okay, now compare that with π × r2 = π × 81 it means r^2 must be 81, in that case, what's the radius?

OpenStudy (anonymous):

it don't say

terenzreignz (terenzreignz):

Of course it doesn't, you can find it... you know that r^2 = 81, so r must be equal to...?

OpenStudy (anonymous):

is the answer 4.5?

OpenStudy (anonymous):

the diameter is twice the radius

OpenStudy (anonymous):

I got it now